Factoring 3^rd Degree Polynomials
Let us solve the polynomial equation, x^3 + x^2 - 20x = 0

Answer by nerdybill(7384)

x^3 + x^2 - 20x = 0
factor out common term among all terms:
x(x^2 + x - 20) = 0
x(x+5)(x-4) = 0
x = {-5, 0, 4}